the measure of angle B, the complement of angle A, is one-sixth the measure of angle C, the supplement of angle A. What are the measures of each of the three angles?
Answer by ewatrrr(24785)


Hi,
Let x represent angle A
B = (90-x)
C= (180-x)
write as we read***
B = C/6
(90-x) = (180-x)/6
solve for x
6(90-x) = 180 -x
540 - 6x = 180 -x
360 = 5x
72 = x
A= 72degrees, B= 18degrees and C= 108degrees
checking our answer
18degrees = 108degrees/6